Tony had an equal number of cranberry bars and walnut bars. He gave away 66 cranberry bars. He then had 4 times as many walnut bars as cranberry bars left. How many bars did he have at first?
Answer by richwmiller(17219)

c=w
4*(c-66)=c
4c-66*4=c
3c=66*4
c=22*4
c=88
w=88